The additive inverse of the number 1 + 4i is:
a. -1 + 4i
b. 1/17 - 4/17 i
c. -1 - 4i
d. 1 - 4i

Answer by venugopalramana(3286)

The additive inverse of the number 1 + 4i is:
additive identity element is 0+0i...since (a+bi)+(0+0i)=a+bi
so additive inverse of 1+4i is -1-4i...since 1+4i-1-4i=0+0i
a. -1 + 4i
b. 1/17 - 4/17 i
c. -1 - 4i
d. 1 - 4i